OpenStudy (anonymous):

Having problem with the fraction part. KE=1/2mv square; solve for v.?

OpenStudy (anonymous):

Firstly multiply by 2 both the sides and then divide by m both the sides..

mathslover (mathslover):

ok so here allow me to correct it : is it : kinetic energy = \(\large{\frac{1}{2}mv^2}\)

OpenStudy (anonymous):

Then take the square root..

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

@mathslover he is right..

mathslover (mathslover):

r u mentioning me that reggie is right?

OpenStudy (anonymous):

Multiply by 2 both the sides first: \[\large mv^2 = 2(K.E.)\] Now divide by m both the sides: \[\large v^2 = \frac{2 \times K.E.}{m}\] Now take the square root both sides: \[\large v = \sqrt{\frac{2 \times K.E.}{m}}\]

OpenStudy (anonymous):

thank you for helping me.